how would you do the equation
3x+21=5x+15?
by doing the subtract from each side for example
z-4=56
3x+21=5x+15
3x+21-21=5x+15-21
3x+0=5x-6
3x-5x=5x-5x-6
-2x=0-6
-2x/-2=-6/-2
x=3
Check:
3x+21 ? 5x+15
3(3)+21 ? 5(3)+15
9+21 ? 15+15
30 = 30 Check, answer is right!
z-4+4=56+4
z=60

Found 2 solutions by checkley77, eperette:
Answer by checkley77(12844)

3x+21=5x+15 subtract 5x & subtract 21
3x-5x=15-21
-2x=-6
x=-6/-2
x=3 answer.

Answer by eperette(173)